# Quantum refraction effects in pulsar radio emission

### v1

@@ 3 Change of polarization of a light ray due to quantum refraction
In Section 2 we have separated Case I and Case II for the two different values of the refractive index $n$ attributed to the same magnetosphere, depending on the propagation and polarization of the light ray associated with them, as given by equation (4), according to the PM Lagrangian model (Kim & Kim (2022)). In accordance with our perturbation analysis, the refractive index can be approximated via expansion as
@@ 3.1 For Case I
According to Born et al. (1999), the propagation of the unit polarization vector $\mathbf{\varepsilon}$ can be described by the equation:
$\frac{\mathrm{d}\mathbf{\varepsilon}}{\mathrm{d}\tau}\equiv n\frac{\mathrm{d}\mathbf{\varepsilon}}{\mathrm{d}s}=-\left(\mathbf{\varepsilon}\cdot\nabla\left(\ln n\right)\right)\nabla\mathcal{S},$ (32)
@@
$\frac{\mathrm{d}\mathbf{\varepsilon}}{\mathrm{d}s}=-\left(\mathbf{\varepsilon}\cdot\nabla\left(\ln n\right)\right)\frac{\mathrm{d}\mathbf{r}}{\mathrm{d}s}.$ (33)
@@
$\frac{\mathrm{d}}{\mathrm{d}s}\left(\delta\mathbf{\varepsilon}_{\left[1\right]}\right)=-\left[\mathbf{\varepsilon}_{\left[0\right]}\cdot\nabla\left(\delta n_{\left[1\right]}\right)\right]\mathbf{\hat{n}}_{\left[0\right]},$ (35)
where $\mathbf{\varepsilon}_{\left[0\right]}$ denotes the classical polarization vector and $\delta\mathbf{\varepsilon}_{\left[1\right]}$ is the leading (first) order quantum correction to it, and $\mathbf{\hat{n}}_{\left[0\right]}$ refers to the classical propagation vector. This equation describes how quantum refraction affects the propagation of our polarization vector along the path of the light ray by means of perturbation.
One possible way of prescribing the polarization vector classically is
$\mathbf{\varepsilon}_{\left[0\right]}=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\hat{n}_{y\left[0\right]}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (36)
[…]
@@
$\mathbf{\varepsilon}_{\mathrm{o}}=\left(1-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{y}-\left(1-\eta_{1}B_{\mathrm{o}}^{2}\right)\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (37)
which has been adapted from its original expression in Kim & Kim (2022) to the geometry of our rotating magnetosphere, with the consideration of equation (36). [^3] Originally, in Kim & Kim (2022) the initial polarization vector associated with $n$ for Case I is given by $\mathbf{\varepsilon}_{\mathrm{o}}=\left(1-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\cos\theta\mathbf{e}_{x}-\left(1-\eta_{1}B_{\mathrm{o}}^{2}\right)\sin\theta\mathbf{e}_{z}$, which corresponds to the classical polarization vector, $\mathbf{\varepsilon}_{\left[0\right]}=$ $\cos\theta\mathbf{e}_{x}-\sin\theta\mathbf{e}_{z}$. Then we may separate the classical part,
$\mathbf{\varepsilon}_{\mathrm{o}\left[0\right]}\equiv\left.\mathbf{\varepsilon}_{\left[0\right]}\right|_{s=0}=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z}$ (38)
@@
$\delta\mathbf{\varepsilon}_{\mathrm{o}\left[1\right]}=\left(-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\hat{n}_{z\left[0\right]}\mathbf{e}_{x}+\eta_{1}B_{\mathrm{o}}^{2}\hat{n}_{x\left[0\right]}\mathbf{e}_{z}.$ (39)
The polarization vector with the first order correction due to the quantum refraction effect can be obtained in a similar manner as in Sect 2. Integrating equation (35) with respect to $s=r_{\mathrm{o}}\lambda$, and combining this with equation (36), and using equations (18), (21), (22), (23), (36) and (39), we finally have $\mathbf{\varepsilon}=\varepsilon_{x}\mathbf{e}_{x}+\varepsilon_{y}\mathbf{e}_{y}+\varepsilon_{z}\mathbf{e}_{z}$ with
@@
$\left.\left.\hskip 8.0pt-\frac{\Omega r_{\mathrm{o}}}{c}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\mathcal{I}\left(\lambda\right)\right.\!\!\!\!\!\right],$ (42)
@@
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (10), (11) and (26), respectively, and $B_{\mathrm{o}}=\mu\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}/r_{\mathrm{o}}^{3}$, and $\mathcal{I}\left(\lambda\right)$ refers to equation (30). In Fig. 3 is plotted the change in the polarization vector $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ against $0\leq\lambda\leq 10$, wherein $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{1}B_{\mathrm{o}}^{2}$, $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$ (fairly exaggerated for the purpose of visualization). From this plot, one can see that the polarization vector changes drastically along the $x$-axis and $z$-axis only at the beginning of the propagation of our light ray while keeping the linear relationship between the components along these axes. Then the $x$- and $z$-components return to the initial values and remain the same for the rest of the light ray propagation.
@@
Figure 3: The change in the polarization vector, $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ plotted against $0\leq\lambda\leq 10$; the red solid curve represents the total polarization vector (classical polarization vector $+$ quantum correction), while the blue and green dashed curves represent the classical polarization vector and the quantum correction, respectively. For the purpose of visualization, the quantum refraction effect is fairly exaggerated by $\eta_{1}B_{\mathrm{o}}^{2}$, $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$, $10^{9}$ times as large as an actual value $\sim 10^{-5}$.
@@
By equations (24)-(26) and (40)-(42) one can inspect the orthogonality between the propagation and polarization vectors,$\mathbf{\ \hat{n}}$ and $\mathbf{\varepsilon}$:
@@
$\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}\approx 2\eta_{2}B_{\mathrm{o}}^{2}\hat{n}_{x\left[0\right]}\hat{n}_{z\left[0\right]}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right),$ (43)
@@
where $\hat{n}_{x\left[0\right]}$ and $\hat{n}_{z\left[0\right]}$ are given by equations (10) and (11), respectively. This implies that the quantum refraction effect results in breaking the orthogonality at the leading order in $\eta_{2}B_{\mathrm{o}}^{2}$. However, the departure from the orthogonality remains constant under this effect, being determined at the leading order in $\eta_{2}B_{\mathrm{o}}^{2}$ solely by the initial conditions for quantum refraction. For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$ and a usual value of $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, we find the departure value to be approximately $3.02\times 10^{-5}\,\mathrm{rad}$.
@@
$\varphi_{\mathrm{F}}\equiv\cos^{-1}\left(\mathbf{\varepsilon}_{\left[0\right]}\cdot\mathbf{\hat{\varepsilon}}\right)\approx\sqrt{\delta\mathbf{\varepsilon}_{\left[1\right]}^{2}-\left(\mathbf{\varepsilon}_{\left[0\right]}\cdot\delta\mathbf{\varepsilon}_{\left[1\right]}\right)^{2}}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right).$ (44)
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$ and a usual value of $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, we find $\varphi_{\mathrm{F}}\approx 3.02\times 10^{-5}\,\mathrm{rad}$.
Note that the values, $\mathbf{\hat{n}}\cdot\mathbf{\varepsilon}\approx 3.02\times 10^{-5}\,\mathrm{rad}$ and $\varphi_{\mathrm{F}}\approx 3.02\times 10^{-5}\,\mathrm{rad}$, obtained as above are small but comparable to our perturbation parameter $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$; therefore, these effects are truly quantum refractive.
@@ 3.2 For Case II
Following Kim & Kim (2022), one can express the initial polarization vector associated with the refractive index $n$ for Case II in accordance with equation (4) as
$\mathbf{\varepsilon}_{\mathrm{o}}=-\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{z},$ (45)
where $\widehat{n}_{x\left[0\right]}$, $\widehat{n}_{z\left[0\right]}$ and $\widehat{n}_{y\left[0\right]}$ are given by equations (10)-(12), respectively. This has been adapted from its original expression in Kim & Kim (2022) to the geometry of our rotating magnetosphere, and at the same time, shall be equal to the classical polarization vector; [^4] Originally, in Kim & Kim (2022) the initial polarization vector associated with $n$ for Case II is given by $\mathbf{\varepsilon}_{\mathrm{o}}=\mathbf{e}_{y}$, which is identical to the classical polarization vector, $\mathbf{\varepsilon}_{\left[0\right]}=$ $\mathbf{e}_{y}$. i.e.,
$\mathbf{\varepsilon}_{\left[0\right]}=\mathbf{\varepsilon}_{\mathrm{o}}.$ (46)
[…]
The polarization vector with the first order correction due to the quantum refraction effect can be obtained in the same manner as in Case I above. That is, we combine the classical polarization vector $\mathbf{\varepsilon}_{\left[0\right]}$ (given by equation (46)) and the quantum correction (given by the integral of equation (35) with respect to $s=r_{\mathrm{o}}\lambda$, wherein the refractive index is expressed with $\eta_{1}$, following equation (31)). We obtain $\mathbf{\varepsilon}=\varepsilon_{x}\mathbf{e}_{x}+\varepsilon_{y}\mathbf{e}_{y}+\varepsilon_{z}\mathbf{e}_{z}$ with
@@
$\!\left.\left.\hskip 10.0pt-\frac{\Omega r_{\mathrm{o}}\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)}{c}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\mathcal{I}\left(\lambda\right)\right.\!\!\!\!\!\right],$ (47)
@@
$\!\left.\left.\hskip 10.0pt+\frac{\Omega r_{\mathrm{o}}\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)}{c}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\mathcal{I}\left(\lambda\right)\right.\!\!\!\!\!\right],$ (48)

### v2

@@ 4 Change of polarization of a light ray due to quantum refraction
In Section 3 we have separated Case I and Case II for the two different values of the refractive index $n$ attributed to the same magnetosphere, depending on the propagation and polarization of the light ray associated with them, as given by equation (9), according to the PM Lagrangian model (Kim & Kim (2022)). In accordance with our perturbation analysis, the refractive index can be approximated via expansion as (Adler (1971))
@@ 4.1 For Case I
According to Born et al. (1999), the propagation of the unit polarization vector $\boldsymbol{\varepsilon}$ can be described by the equation:
$\frac{\mathrm{d}\boldsymbol{\varepsilon}}{\mathrm{d}\tau}\equiv n\frac{\mathrm{d}\boldsymbol{\varepsilon}}{\mathrm{d}s}=-\left(\boldsymbol{\varepsilon}\cdot\nabla\left(\ln n\right)\right)\nabla\mathcal{S},$ (37)
@@
$\frac{\mathrm{d}\boldsymbol{\varepsilon}}{\mathrm{d}s}=-\left(\boldsymbol{\varepsilon}\cdot\nabla\left(\ln n\right)\right)\frac{\mathrm{d}\mathbf{r}}{\mathrm{d}s}.$ (38)
@@
$\frac{\mathrm{d}}{\mathrm{d}s}\left(\delta\boldsymbol{\varepsilon}_{\left[1\right]}\right)=-\left[\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\nabla\left(\delta n_{\left[1\right]}\right)\right]\mathbf{\hat{n}}_{\left[0\right]},$ (40)
where $\boldsymbol{\varepsilon}_{\left[0\right]}$ denotes the classical polarization vector and $\delta\boldsymbol{\varepsilon}_{\left[1\right]}$ is the leading (first) order quantum correction to it, and $\mathbf{\hat{n}}_{\left[0\right]}$ refers to the classical propagation vector. This equation describes how quantum refraction affects the propagation of our polarization vector along the path of the light ray by means of perturbation.
One possible way of prescribing the polarization vector classically, with the consideration of the rotational effect is
$\boldsymbol{\varepsilon}_{\left[0\right]}=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\hat{n}_{y\left[0\right]}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (41)
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (15)-(17), respectively. It can be easily checked out that $\boldsymbol{\varepsilon}_{\left[0\right]}$ is orthogonal to the propagation vector, $\mathbf{\hat{n}}_{\left[0\right]}=\hat{n}_{x\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]}\mathbf{e}_{z}$, i.e., $\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\mathbf{\hat{n}}_{\left[0\right]}\approx 0+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2}\right)$ while it is normalized, i.e., $\boldsymbol{\varepsilon}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2}\right)$.
@@
$\boldsymbol{\varepsilon}_{\mathrm{o}}=\left(1-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{y}-\left(1-\eta_{1}B_{\mathrm{o}}^{2}\right)\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (42)
which has been adapted from its original expression in Kim & Kim (2022) to the geometry of our rotating magnetosphere, with the consideration of equation (41). [^4] Originally, in Kim & Kim (2022) the initial polarization vector associated with $n$ for Case I is given by $\boldsymbol{\varepsilon}_{\mathrm{o}}=\left(1-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\cos\theta\mathbf{e}_{x}-\left(1-\eta_{1}B_{\mathrm{o}}^{2}\right)\sin\theta\mathbf{e}_{z}$, which corresponds to the classical polarization vector, $\boldsymbol{\varepsilon}_{\left[0\right]}=$ $\cos\theta\mathbf{e}_{x}-\sin\theta\mathbf{e}_{z}$. Then we may separate the classical part,
$\boldsymbol{\varepsilon}_{\mathrm{o}\left[0\right]}\equiv\left.\boldsymbol{\varepsilon}_{\left[0\right]}\right|_{s=0}=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}\pm\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z}$ (43)
@@
$\delta\boldsymbol{\varepsilon}_{\mathrm{o}\left[1\right]}=\left(-\eta_{1}B_{\mathrm{o}}^{2}+2\eta_{2}B_{\mathrm{o}}^{2}\right)\hat{n}_{z\left[0\right]}\mathbf{e}_{x}+\eta_{1}B_{\mathrm{o}}^{2}\hat{n}_{x\left[0\right]}\mathbf{e}_{z}.$ (44)
The polarization vector with the first order correction due to the quantum refraction effect can be obtained in a similar manner as in Sect 3. Integrating equation (40) with respect to $s=r_{\mathrm{o}}\lambda$, and combining this with equation (41), and using equations (23), (26), (27), (28), (41) and (44), we finally have $\boldsymbol{\varepsilon}=\varepsilon_{x}\mathbf{e}_{x}+\varepsilon_{y}\mathbf{e}_{y}+\varepsilon_{z}\mathbf{e}_{z}$ with
@@
$\left.\left.\hskip 8.0pt-\frac{\Omega r_{\mathrm{o}}\sin\alpha}{c\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\mathcal{I}\left(\lambda\right)\right._{\,{}_{\,{}_{\,{}_{\,}}}}^{\,{}^{\,{}^{\,{}^{\,}}}}\!\!\!\!\!\right],$ (47)
@@
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (15), (16) and (31), respectively, and $B_{\mathrm{o}}=\mu\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}/r_{\mathrm{o}}^{3}$, and $\mathcal{I}\left(\lambda\right)$ refers to equation (35). In Fig. 3 is plotted the change in the polarization vector $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ against $0\leq\lambda\leq 10$ on a logarithmic scale, wherein $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$, $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$ and $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$. From this plot, one can see that the total polarization vector changes drastically along the $x$-axis and $z$-axis near the beginning of the propagation of our light ray due to the quantum refraction effect, while the classical polarization vector changes only along the $y$-axis due to the rotational effect, as can be seen from equation (41).
@@
Figure 3: The change in the polarization vector, $\left(\Delta\varepsilon_{x},\Delta\varepsilon_{y},\Delta\varepsilon_{z}\right)\equiv\left.\left(\varepsilon_{x}\left(\lambda\right),\varepsilon_{y}\left(\lambda\right),\varepsilon_{z}\left(\lambda\right)\right)\right|_{0}^{\lambda}$ plotted against $0\leq\lambda\leq 10$ on a logarithmic scale; the red solid curve and the blue dashed curve represent the total polarization vector (classical polarization vector $+$ quantum correction) and the classical polarization vector, respectively.
@@
By equations (29)-(31) and (45)-(47) one can inspect the orthogonality between the propagation and polarization vectors,$\mathbf{\ \hat{n}}$ and $\boldsymbol{\varepsilon}$:
@@
$\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}\approx 2\eta_{2}B_{\mathrm{o}}^{2}\hat{n}_{x\left[0\right]}\hat{n}_{z\left[0\right]}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right),$ (48)
@@
where $\hat{n}_{x\left[0\right]}$ and $\hat{n}_{z\left[0\right]}$ are given by equations (15) and (16), respectively. This implies that the quantum refraction effect results in breaking the orthogonality at the leading order in $\eta_{2}B_{\mathrm{o}}^{2}$. However, the departure from the orthogonality remains constant under this effect, being determined at the leading order in $\eta_{2}B_{\mathrm{o}}^{2}$ solely by the initial conditions for quantum refraction. For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$ and a usual value of $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, we find the departure value to be $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}\right)\approx 3.02\times 10^{-5}\,\mathrm{rad}$.
@@
$\varphi_{\mathrm{F}}\equiv\cos^{-1}\left(\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\boldsymbol{\hat{\varepsilon}}\right)\approx\sqrt{\delta\boldsymbol{\varepsilon}_{\left[1\right]}^{2}-\left(\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\delta\boldsymbol{\varepsilon}_{\left[1\right]}\right)^{2}}+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2},_{\left[2\right]}\right).$ (49)
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$ and usual values of $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$, $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, we find $\varphi_{\mathrm{F}}\approx 3.02\times 10^{-5}\,\mathrm{rad}$.
Note that the values, $\sin^{-1}\left(\mathbf{\hat{n}}\cdot\boldsymbol{\varepsilon}\right)\approx 3.02\times 10^{-5}\,\mathrm{rad}$ and $\varphi_{\mathrm{F}}\approx 3.02\times 10^{-5}\,\mathrm{rad}$, obtained as above are small but comparable to our perturbation parameters, $\eta_{1}B_{\mathrm{o}}^{2}\approx 2.45\times 10^{-5}$ and $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$; therefore, these effects are truly quantum refractive.
@@ 4.2 For Case II
One can write down the classical polarization vector associated with the refractive index $n$ for Case II as given by (9), with the consideration of the rotational effect as
$\boldsymbol{\varepsilon}_{\left[0\right]}=-\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{z},$ (50)
where $\hat{n}_{x\left[0\right]}$, $\hat{n}_{z\left[0\right]}$ and $\hat{n}_{y\left[0\right]}$ are given by equations (15)-(17), respectively. In accordance with this, the initial polarization vector shall be
$\boldsymbol{\varepsilon}_{\mathrm{o}}=-\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\left.\hat{n}_{y\left[0\right]}\right|_{s=0}\mathbf{e}_{z},$ (51)
which has been adapted from its original expression in Kim & Kim (2022) to the geometry of our rotating magnetosphere. [^5] In Kim & Kim (2022) the initial polarization vector associated with $n$ for Case II is given by $\boldsymbol{\varepsilon}_{\mathrm{o}}=\mathbf{e}_{y}$, which is identical to the classical polarization vector, $\boldsymbol{\varepsilon}_{\left[0\right]}=\mathbf{e}_{y}$.
It can be checked out that $\boldsymbol{\varepsilon}_{\left[0\right]}$ is orthogonal to the propagation vector, $\mathbf{\hat{n}}_{\left[0\right]}=\hat{n}_{x\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]}\mathbf{e}_{z}$, i.e., $\boldsymbol{\varepsilon}_{\left[0\right]}\cdot\mathbf{\hat{n}}_{\left[0\right]}=0$ while it is normalized, i.e., $\boldsymbol{\varepsilon}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega r_{\mathrm{o}}/c\right)^{2}\right)$. In addition, from equations (41) and (50) we find that $\boldsymbol{\varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}I}\right)}\cdot\boldsymbol{\varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}II}\right)}=0$. Then one can note that the three vectors, $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\ \varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}I}\right)}$ and $\boldsymbol{\varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}II}\right)}$ form a classical orthonormal basis. [^6] From Appendix A one can see that $\mathbf{\hat{n}}_{\left[0\right]}=\mathbf{T}_{\left[0\right]}=\hat{n}_{x\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]}\mathbf{e}_{z}$, $\boldsymbol{\varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}I}\right)}=\mathbf{N}_{\left[0\right]}=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z}$ and $\boldsymbol{\varepsilon}_{\left[0\right]\,\left(\mathrm{Case~{}II}\right)}=\mathbf{B}_{\left[0\right]}=-\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{z}$, where $\mathbf{T}_{\left[0\right]}$, $\mathbf{N}_{\left[0\right]}$ and $\mathbf{B}_{\left[0\right]}$ denote the unperturbed (classical) part of the unit tangent, normal and bi-normal vectors, taken from equations (60)-(62), respectively, with all the terms of $\delta n_{\left[1\right]}$ removed.
The polarization vector with the first order correction due to the quantum refraction effect can be obtained in the same manner as in Case I above. That is, we combine the classical polarization vector $\boldsymbol{\varepsilon}_{\left[0\right]}$ (given by equation (50)) and the quantum correction (given by the integral of equation (40) with respect to $s=r_{\mathrm{o}}\lambda$, wherein the refractive index is expressed with $\eta_{1}$, following equation (36)). We obtain $\boldsymbol{\varepsilon}=\varepsilon_{x}\mathbf{e}_{x}+\varepsilon_{y}\mathbf{e}_{y}+\varepsilon_{z}\mathbf{e}_{z}$ with
@@
$\!\left.\left.\hskip 10.0pt-\frac{\Omega r_{\mathrm{o}}\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\sin\alpha}{c\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\mathcal{I}\left(\lambda\right)\right._{\,{}_{\,{}_{\,{}_{\,}}}}^{\,{}^{\,{}^{\,{}^{\,}}}}\!\!\!\!\!\right],$ (52)
@@
$\!\left.\left.\hskip 10.0pt+\frac{\Omega r_{\mathrm{o}}\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\sin\alpha}{c\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\mathcal{I}\left(\lambda\right)\right._{\,{}_{\,{}_{\,{}_{\,}}}}^{\,{}^{\,{}^{\,{}^{\,}}}}\!\!\!\!\!\right],$ (53)